OpenStudy (anonymous):

Does the integral of pi/(x^2) with the limits of [1, inf] converge?

OpenStudy (anonymous):

So this is what I get, but my calculator says that it actually diverges: \[\int\limits_{1}^{\infty} \Pi/(x^2)= \lim_{1 \rightarrow \infty}-\Pi/x=0-\Pi(-1)=\Pi\]